iosangel:

Factor the expression completely over the complex numbers. y^4+14y^2+49

imbadkidjenna:

1. y^2x2 + 2 x y^2 x 7 + 7^2 2. ( y^2 )^2 + 2 x 7 + 7^2 3. (y^2 + 7)^2

iosangel:

Thats not right

Vocaloid:

sorry this took so long (y^2 + 7)^2 = 0 taking the square root of both sides y^2 + 7 = 0 y^2 = -7 y = +/- i * sqrt(7) so we can re-write y^2 + 7 = 0 as (y + i *sqrt(7))(y - i * sqrt(7)) going back to the original polynomial, (y^2 + 7)^2 is just the square of the previous result, so simply square both factors (y + i *sqrt(7))^2 * (y - i * sqrt(7))^2 = (y + i *sqrt(7))*(y + i *sqrt(7))*(y - i *sqrt(7))*(y - i *sqrt(7))